# Localization of information driven by stochastic resetting

### v1

@@ Appendix C Case of a parabolic VDLE
$\lambda(v)=\lambda_{0}\left[1-\left(\frac{v}{v_{\rm B}}\right)^{2}\right],$ (42)
@@
$\tilde{\lambda}(v)=\lambda_{0}\left[1-\left(\frac{v}{v_{\rm B}}\right)^{2}\right]-r,$ (43)
@@
$\tilde{\Lambda}(k)=\lambda_{0}\left[1-\left(\frac{v_{\rm B}}{2\lambda_{0}}k\right)^{2}\right]-r$ (44)
@@
$\tilde{\lambda}(v<v^{*})=-2\frac{v}{v_{\rm B}}\sqrt{\lambda_{0}(r-r_{\rm c})}$ (45)
@@
$v^{*}=v_{\rm B}\sqrt{\frac{r-r_{\rm c}}{\lambda_{0}}}.$ (46)
The expression of $\tilde{\lambda}(v\geq v^{*})$ is simply given by Eq. (43) and $\tilde{v}_{\rm B}=0$. The Lyapunov spectrum is computed using Eq. (6),
@@
$\quad+\left.\int_{v^{*}}^{\infty}\!\!{\rm d}v\,\cos(kvt)\,{\rm e}^{\left(\lambda_{0}\left[1-(v/v_{\rm B})^{2}\right]-r\right)t}\right|.$ (47)
@@
$\tilde{D}_{\rm st}(x,x^{\prime})=r\int_{0}^{\infty}{\rm d}\tau\,{\rm e}^{\left[\lambda\left(\frac{x-x^{\prime}}{\tau}\right)-r\right]\tau}.$ (49)
@@
$\tilde{D}_{\rm st}(x,x^{\prime})=\frac{r}{r-r_{\rm c}}\frac{|x-x^{\prime}|}{\xi_{r}}\ K_{1}\left(|x-x^{\prime}|/\xi_{r}\right),$ (50)
where $K_{1}(x)$ is a modified Bessel function of the second kind. At large distances, using $K_{1}(x)\sim\sqrt{\pi/2x}\exp(-x)$, the expression above describes an OTOC that is exponentially localized around the perturbation site with a localization length $\xi_{r}=v_{\rm B}/\sqrt{4\lambda_{0}(r-r_{\rm c})}$ that diverges at the dynamical transition with an exponent $\nu=1/2$.
@@ Appendix D Case of a discrete chain
The case of a parabolic VDLE can also be explicitly solved on an infinite one-dimensional chain. We consider the parabolic velocity-dependent Lyapunov given in Eq. (42) and set the lattice spacing to unity. The eigenvalues of the OTOC are obtained by discrete Fourier transforms, yielding
$D(k;t)={\rm e}^{\lambda_{0}t}\,\vartheta\left(\frac{k}{2},\exp\left(-\frac{\lambda_{0}}{v_{\rm B}^{2}t}\right)\right),$ (51)
@@
$D(k;t)\sim\sqrt{\pi\frac{v_{{\rm B}}^{2}t}{\lambda_{0}}}\;{\rm e}^{\lambda_{0}\left[1-\left(\frac{v_{{\rm B}}}{2\lambda_{0}}k\right)^{2}\right]t},$ (52)
@@
$\Lambda(k)=\lambda_{0}\left[1-\left(\frac{v_{\rm B}}{2\lambda_{0}}k\right)^{2}\right],$ (53)

### v2

@@ Appendix C Case of a parabolic VDLE
$\lambda(v)=\lambda_{0}\left[1-\left(\frac{v}{v_{\rm B}}\right)^{2}\right],$ (A23)
@@
$\tilde{\lambda}(v)=\lambda_{0}\left[1-\left(\frac{v}{v_{\rm B}}\right)^{2}\right]-r,$ (A24)
@@
$\tilde{\Lambda}(k)=\lambda_{0}\left[1-\left(\frac{v_{\rm B}}{2\lambda_{0}}k\right)^{2}\right]-r$ (A25)
@@
$\tilde{\lambda}(v<v^{*})=-2\frac{v}{v_{\rm B}}\sqrt{\lambda_{0}(r-r_{\rm c})}$ (A26)
@@
$v^{*}=v_{\rm B}\sqrt{\frac{r-r_{\rm c}}{\lambda_{0}}}.$ (A27)
The expression of $\tilde{\lambda}(v\geq v^{*})$ is simply given by Eq. (A24) and $\tilde{v}_{\rm B}=0$. The Lyapunov spectrum is computed using Eq. (6),
@@
$\quad+\left.\int_{v^{*}}^{\infty}\!\!{\rm d}v\,\cos(kvt)\,{\rm e}^{\left(\lambda_{0}\left[1-(v/v_{\rm B})^{2}\right]-r\right)t}\right|.$ (A28)
@@
$\tilde{D}_{\rm st}(x,x^{\prime})=r\int_{0}^{\infty}{\rm d}\tau\,{\rm e}^{\left[\lambda\left(\frac{x-x^{\prime}}{\tau}\right)-r\right]\tau}.$ (A30)
@@
$\tilde{D}_{\rm st}(x,x^{\prime})=\frac{r}{r-r_{\rm c}}\frac{|x-x^{\prime}|}{\xi_{r}}\ K_{1}\left(|x-x^{\prime}|/\xi_{r}\right),$ (A31)
where $K_{1}(x)$ is a modified Bessel function of the second kind. At large distances, using $K_{1}(x)\sim\sqrt{\pi/2x}\exp(-x)$, the expression above describes an OTOC that is exponentially localized around the perturbation site with a localization length $\xi_{r}=v_{\rm B}/\sqrt{4\lambda_{0}(r-r_{\rm c})}$ that diverges at the dynamical transition with an exponent $\nu=1/2$. At late times, the approach to the steady state follows
$\tilde{D}_{\rm st}(x,x^{\prime})-\tilde{D}(x,x^{\prime};t)\simeq{\rm e}^{(\lambda_{0}-r)t}\frac{\lambda_{0}}{r-\lambda_{0}}.$ (A32)
This exhibits a timescale $\tau_{r}:=1/(r-r_{\rm c})$ that also diverges at the dynamical transition. Altogether, this critical phenomenon is controlled by a dynamical exponent $z=2$.
@@ Appendix D Case of a discrete chain
The case of a parabolic VDLE can also be explicitly solved on an infinite one-dimensional chain. We consider the parabolic velocity-dependent Lyapunov given in Eq. (A23) and set the lattice spacing to unity. The eigenvalues of the OTOC are obtained by discrete Fourier transforms, yielding
$D(k;t)={\rm e}^{\lambda_{0}t}\,\vartheta\left(\frac{k}{2},\exp\left(-\frac{\lambda_{0}}{v_{\rm B}^{2}t}\right)\right),$ (A33)
@@
$D(k;t)\sim\sqrt{\pi\frac{v_{{\rm B}}^{2}t}{\lambda_{0}}}\;{\rm e}^{\lambda_{0}\left[1-\left(\frac{v_{{\rm B}}}{2\lambda_{0}}k\right)^{2}\right]t},$ (A34)
@@
$\Lambda(k)=\lambda_{0}\left[1-\left(\frac{v_{\rm B}}{2\lambda_{0}}k\right)^{2}\right],$ (A35)